OpenStudy (anonymous):

ALGEBRA HELP PLEASE!!!!!! WILL FAN & MEDAL :) 1. Simplify the radical expression. square root of 45 A) 5 square root 3 B) 3 square root 5 C) 3 square root 15 D) 15 2. Simplify the radical expression. square root of 56x^2 A) 28x B) 2x square root 14 C) 2x square root 7 D) 2 square root 14x^2

Mehek (mehek14):

\(\sqrt{45}\) factors of 45?

OpenStudy (calculusxy):

\(\large \sqrt{45}\) first list the factors of 45

Mehek (mehek14):

\(\sqrt{9}=\)

OpenStudy (calculusxy):

good now do you notice a perfect square from either one of those factors?

OpenStudy (anonymous):

1,3,5,9,15,45

OpenStudy (calculusxy):

we can work with 5 and 9

OpenStudy (anonymous):

Square root of 9 = 3

Mehek (mehek14):

so 3 goes out side and 5 stays inside \(3\sqrt{5}\)

OpenStudy (calculusxy):

i will let @Mehek14 help you with the first one. seems like @Mehek14 got that question.

Mehek (mehek14):

so simplified is \(3\sqrt{5}\)

Mehek (mehek14):

I actually have to go :(

OpenStudy (calculusxy):

i will take over now if you may :)

OpenStudy (calculusxy):

so the second question is almost like the first one. \(\sqrt{56x^2}\) this is basically saying \(\sqrt{56 \times x^2}\) which means we need to find the square root of 56 and \(x^2\) individually.

OpenStudy (calculusxy):

so list the factors of 56 and make sure that one of the factors (of the two) is a perfect square (if there is any)

OpenStudy (anonymous):

@calculusxy Factors of 56: 1, 2, 4, 7, 14, 28

OpenStudy (calculusxy):

but 1 x 2 is not 28

OpenStudy (calculusxy):

but if we have 4 and 7 (4 x 7 = 28), then we have 4 as the perfect square.

OpenStudy (calculusxy):

2 is not a perfect square because it leads to 1.414...

OpenStudy (calculusxy):

we will stick with 4 and 7. do you understand why?

OpenStudy (calculusxy):

@Answers101

OpenStudy (calculusxy):

okay so \(\sqrt{4} = ?\)

OpenStudy (calculusxy):

good so we will place 2 out of the parenthesis and have 7 inside of the parenthesis (because it is not a perfect square) we now have to find the square root of \(x^2\). what will that be?

OpenStudy (calculusxy):

not really because x^2 is the same thing as \(x \times x\) so the square root of x^2 = x. do you get it?

OpenStudy (anonymous):

Can you help me with a few more? I can open a new question

OpenStudy (calculusxy):

okay so that will be placed outside of the square root as well now we have \(\large 2x \sqrt{7}\)

OpenStudy (calculusxy):

i can only help with one more b/c i have to go soon too